d/dx[(x^2+1)^sinx] .
1
Expert's answer
2020-09-16T19:16:53-0400

Solution:

To find "\frac{d}{dx} [(x^2+1)^{\sin x }]"

Let "y=(x^2+1)^{\sin x } ...(i)"

Taking log to both sides,

"\log y=\log (x^2+1)^{\sin x }
\\ \log y=\sin x \log (x^2+1) [\because \log m^n=n \log m]"

Now differentiating both sides w.r.t "x" using product rule on right side,

"\frac d {dx} \log y=\sin x \frac d {dx} [\log (x^2+1)]+\log (x^2+1) \frac d {dx}\sin x"

"\frac 1 y\frac {dy} {dx}=\sin x[\frac 1 {x^2+1} \frac d {dx} (x^2+1)]+\log (x^2+1) \cos x
\\ \frac 1 y\frac {dy} {dx}=\sin x[\frac 1 {x^2+1}\times 2x]+\log (x^2+1) \cos x
\\ \frac {dy} {dx}=y[\frac {2x\sin x} {x^2+1}+\log (x^2+1) \cos x]"

Now using (i),

"\frac {dy}{dx}=(x^2+1)^{\sin x}[\frac {2x\sin x} {x^2+1}+\log (x^2+1) \cos x]"

Answer:

"(x^2+1)^{\sin x}[\frac {2x\sin x} {x^2+1}+\log (x^2+1) \cos x]"